THE LOWEST^{3}(n, π*) STATE OF 2-CYCLOPENTEN-1-ONE: CAVITY RINGDOWN ABSORPTION SPEC-TRUM AND RING-BENDING POTENTIAL ENERGY FUNCTION

NATHAN R. PILLSBURY, \underline{STEPHEN DRUCKER}, Department of Chemistry, University of Wisconsin-Eau Claire, Eau Claire, WI 54702; JAEBUM CHOO, Department of Chemistry, Hanyang University, Ansan 425-791, Korea; JAAN LAANE, Department of Chemistry, Texas A & M University, College Station, TX 77843.

The room-temperature cavity ringdown absorption spectra of 2-cyclopenten-1-one (2CP) and deuterated derivatives were recorded near 385 nm. The very weak (ϵ < 1 M^{−1} cm^{−1}) band system in this region is due to the T ← S_{0} electronic transition, where T is the lowest-energy^{3}(n, π*) state. The origin band was observed at 25,963.6 cm^{−1} for the undeuterated molecule and at 25,959.4 and 25,956.2 cm^{−1}, respectively, for 2CP-5-d_{1} and 2CP-5,5-d_{2}. For the −d_{0} isotopomer, about 50 vibronic transitions have been assigned in a region from −500 to +500 cm^{−1} relative to the origin band. Nearly every corresponding assignment was made in the −d_{2} spectrum. Several excited-state fundamentals have been determined for the d_{0}/d_{2} isotopomers, including ring-twisting (ν′_{29} = 238.9/227.8 cm^{−1}), out-of-plane carbonyl deformation (ν′_{28} = 431.8/420.3 cm^{−1}), and in-plane carbonyl deformation (ν′_{19} = 346.3/330.2 cm^{−1}). The ring-bending (ν′_{30}) levels for the T state were determined to be at 36.5, 118.9, 213.7, 324.5, and 446.4 cm^{−1} for the undeuterated molecule. These drop to 29.7, 101.9, 184.8, 280.5, and 385.6 cm^{−1} for the −d_{2} molecule. A potential energy function of the form V = ax^{4} +bx^{2} was fit to the ring-bending levels for each isotopic species. The fitting procedure utilized a kinetic energy expansion that was calculated based on the structure obtained for the T state from ab initio calculations.The barrier to planarity, determined from the best-fitting potential energy functions for the −d_{0}, −d_{1}, and −d_{2} species, ranges from 42.0 to 43.5 cm^{−1}. In the T state, electron repulsion resulting from the spin flip favors nonplanarity. The S_{0} and S_{1} states have planar structures that are stabilized by conjugation.